# GeoSurDepth: Spatial Geometry-Consistent Self-Supervised Depth Estimation for Surround-View Cameras

### v1

@@ Abstract
Accurate surround-view depth estimation provides a competitive alternative to laser-based sensors and is essential for 3D scene understanding in autonomous driving. While prior studies have proposed various approaches that primarily focus on enforcing cross-view constraints at the photometric level, few explicitly exploit the rich geometric structure inherent in both monocular and surround-view setting. In this work, we propose GeoSurDepth, a framework that leverages geometry consistency as the primary cue for surround-view depth estimation. Concretely, we utilize foundation models as a pseudo geometry prior and feature representation enhancement tool to guide the network to maintain surface normal consistency in spatial 3D space and regularize object- and texture-consistent depth estimation in 2D. In addition, we introduce a novel view synthesis pipeline where 2D-3D lifting is achieved with dense depth reconstructed via spatial warping, encouraging additional photometric supervision across temporal, spatial, and spatial-temporal contexts, and compensating for the limitations of single-view image reconstruction. Finally, a newly-proposed adaptive joint motion learning strategy enables the network to adaptively emphasize informative spatial geometry cues for improved motion reasoning. Extensive experiments on DDAD and nuScenes demonstrate that GeoSurDepth achieves state-of-the-art performance, validating the effectiveness of our approach. Our framework highlights the importance of exploiting geometry coherence and consistency for robust self-supervised multi-view depth estimation.
@@ I Introduction
Depth estimation is a fundamental task for 3D scene understanding in autonomous driving. In recent years, self-supervised monocular depth estimation has emerged as a promising approach for 3D perception, eliminating the need for dense groundtruth annotations and making vision-based solutions attractive for large-scale, low-cost deployment [1]. By leveraging photometric reconstruction between consecutive frames or stereo pairs [2][3], these methods can learn depth directly from raw image sequences. Classical self-supervised approaches, particularly those based on monocular video, typically enforce photometric and smoothness constraints to regularize depth estimations through structure-from-motion (SfM). While effective in single-view scenarios, these methods often suffer from scale ambiguity, temporal inconsistency, and limited geometric reasoning, especially in complex scenes with dynamic objects or texture-less regions.
@@
Recently, surround-view depth estimation has received growing attention in autonomous driving and robotics, where multiple cameras collectively capture a $360^{\circ}$ field of view (FoV) [4]. Accurate depth estimation in this setting is crucial for robust scene understanding, obstacle avoidance, and multi-camera fusion. However, extending self-supervised depth estimation to surround-view setups introduces additional challenges, including cross-view consistency, occlusion handling, and spatial alignment across cameras. Although several empirical studies have begun addressing these challenges, existing methods still fall short of fully exploiting the rich geometric relationships present in overlapping views, and often fail to properly disentangle photometry, motion, and spatial geometry for consistent multi-camera depth estimation.
In this work, we propose GeoSurDepth, a framework designed to address the challenges of geometry consistency in surround-view depth estimation through simple yet effective strategies. Our key contributions are summarized as follows: (1) We leverage the powerful depth estimation model DepthAnything V2 (DA) as an indirect pseudo-prior for spatial geometry guidance in surround-view settings. This facilitates accurate and edge-aware self-supervised depth estimation by enforcing 3D surface normal consistency and regularizing object-level and texture-level consistency in 2D. Furthermore, we introduce a cross-modal attention module based on CLIP within the depth network to enhance geometric and semantic feature representation. (2) We propose a novel image synthesis approach, where dense depth reconstructed via spatial warping is utilized to achieve 2D-3D lifting, enabling photometric supervision across temporal, spatial, and spatial-temporal domains. This also provides a complementary supervision signal, compensating for the limitations of image reconstruction using depth estimated in the target view only. (3) An adaptive joint motion learning strategy is introduced to enhance the network’s interpretability in emphasizing informative camera views for motion cues and learning.
In general, in this work, we aim to fully exploit geometry consistency as priors or cues to facilitate surround-view depth estimation by tailoring loss function with geometric priors and features, adapting geometry-driven motion learning and enhancing feature representation from geometric perspective.
@@ II Related Works
FSM [4] is the first work to introduce self-supervised depth estimation to the surround-view setting, aiming to achieve omni-directional dense depth perception. By additionally incorporating photometric reconstruction losses in spatial and spatial-temporal contexts, together with a multi-camera pose consistency constraint, FSM enables scale-aware metric depth estimation by explicitly exploiting spatial geometry across views. To further constrain motion estimation and enhance cross-view interaction, SurroundDepth [5] estimates a single joint vehicle motion rather than independent motions for each camera, and employs a Cross-View Transformer to enrich multi-view feature representations. Subsequently, VFDepth [6] adopts a unified volumetric feature fusion strategy, enabling depth estimation from arbitrary viewpoints. In addition, it proposes a canonical motion estimation strategy that provides a global constraint for the surround-view system and derives per-camera motions via extrinsics-based motion distribution. MCDP [7] leverages the output of a pre-trained DepthAnything V1 model [8] as pseudo-depth for conditional denoising learning. By integrating and iteratively refining cross-view features as conditional inputs, MCDP further improves depth estimation performance. Despite these advances, existing methods primarily focus on pose consistency, feature fusion, or depth refinement, while the explicit geometric consistency and constraints of depth or motion estimation across surround views remains underexplored, limiting their ability to fully exploit the structural relationships inherent in surround-view camera systems.
@@ III-A Problem Formulation
Figure 2: Network architecture of GeoSurDepth. Outputs of DA serve as surround-view geometry priors. Surround-view images at the target frame are first resized to $(518,518)$ before being fed into DA, and the output are interpolated back to original resolution. (a) Adaptive joint motion learning; (b) Cross-modal attention mechanism: For CLIP model, input images are resized to $(214,214)$ for token extraction.
We formulate surround-view depth estimation in a self-supervised manner under the conventional SfM paradigm, where dense depth and ego-motion are jointly learned from multi-camera image sequences. An overview of the proposed architecture is shown in Fig.2. The framework consists of a trainable depth network and pose network, together with frozen foundation models including DepthAnything V2 [9] and CLIP [10] model.
Given surround-view images $\{\mathbf{I}_{i}^{t}\}_{i=1}^{N}$ captured by $N$ cameras at time $t$, a depth encoder extracts multi-view features, which are jointly enhanced by fusing CLIP outputs via a cross-modal attention mechanism to improve geometric–semantic coherence. The enhanced features are then passed to a depth decoder to produce surround-view depth estimates $\{\hat{\mathbf{D}}_{i}^{t}\}_{i=1}^{N}$, which are used for 3D reconstruction and view synthesis within an SfM-based framework. To provide explicit geometric guidance, images at target time are also fed into DA, whose outputs serve as pseudo geometry priors that guide depth network toward geometry-consistent estimations.
@@ III-B Spatial Geometry Priors-guided Self-supervised Training
$\mathbf{p}_{ij}^{t\rightarrow t^{\prime}}=\mathbf{\Pi}_{ij}^{t\rightarrow t^{\prime}}\mathbf{p}_{i}^{t},~\tilde{\mathbf{I}}_{ij}^{t\rightarrow t^{\prime}}(\mathbf{p})=\mathbf{I}_{j}^{t^{\prime}}\left<\mathbf{p}_{ij}^{t\rightarrow t^{\prime}}\right$>, (1)
@@
$\tilde{\mathbf{D}}_{j}(\mathbf{p})=(\tilde{\mathbf{P}}_{j})_{z}\left<\mathbf{p}_{i\rightarrow j}\right$>, (8)
where $\mathbf{P}_{j}\in\mathbb{R}^{3}$ denotes a 3D point in coordinate frame of camera $j$. $(\cdot)_{z}$ implies $z$ value of a point cloud.
@@
$\mathcal{L}_{\text{SDC}}=\sum_{j\in\mathcal{A}(i)}\|\mathbf{D}_{i}-\tilde{\mathbf{D}}_{j}\|_{1},$ (9)
where $\mathcal{A}(i)$ indicates adjacent view of camera $i$.
@@
Geometric prior-guided surface normal consistency (SNC) loss. Foundation models, trained on large-scale datasets with strong generalization capability, can provide valuable prior information for network training. MCDP [7] leverages depth estimations from a pre-trained DepthAnything V1 model [8] as pseudo-depth for conditional denoising learning. VFM-Depth [14] incorporates DINOv2 into the depth encoder to provide universal and stable semantic features. Instead of directly using depth outputs from foundation models as priors, we compute the corresponding surface normal vector maps and enforce 3D consistency in a scale- and shift-invariant manner. We present the theoretical derivation as follows.
Concretely, the output of DA is first normalized to the range $[0,1]$ to mitigate shift effects, yielding a disparity-like representation (see Fig.3). This representation is then converted into a pseudo depth map by applying a clipping range based on the estimated depth (refer to Supplementary 1.1 for more details). Notably, although this procedure produces a depth map, it remains scale-ambiguous. Therefore, we do not use it for direct depth supervision. Instead, we exploit this scale-ambiguous pseudo depth to compute surface normal vector map, which is inherently invariant to scale ambiguity. Specifically, we assume $D^{\text{true}}\approx\gamma D^{\text{DA}}$, where $\gamma$ denotes an unknown scale factor. To obtain surface normal vector, we lift each pixel $\mathbf{p}$ to 3D space as $\mathbf{P}=D^{\text{true}}\mathbf{K}^{-1}\mathbf{p}$. Likewise, we lift two neighboring pixels $\mathbf{p}_{1}$ and $\mathbf{p}_{2}$, chosen such that $\cos(\overrightarrow{\mathbf{p}\mathbf{p}_{1}},\overrightarrow{\mathbf{p}\mathbf{p}_{2}})=1$. The surface normal vector at $\mathbf{P}$ can be computed via cross-product as,
@@
Specifically, we consider eight neighbors of a pixel $\mathbf{p}$ and construct eight ordered pixel pairs $\mathcal{P}(\mathbf{p})=\{(\mathbf{p}_{j_{0}},\mathbf{p}_{j_{1}})\}_{j=1}^{8}$ whose offset vectors relative to $\mathbf{p}$ are mutually perpendicular and arranged in a counterclockwise order. Following (2), we lift these pixels using depth estimation by the depth network and the scale-ambiguous depth of DA, yielding the corresponding 3D point pairs $\mathcal{P}(\mathbf{P})=\{\mathbf{P}_{j_{0}},\mathbf{P}_{j_{1}}\}_{j=1}^{8}$. For each pair, we compute their cross product as described above, from which the surface normal vector map $\mathbf{N}$ is obtained as,
$\mathbf{n}_{j}(\mathbf{P})=\frac{\overrightarrow{\mathbf{P}\mathbf{P}_{j_{0}}}\times\overrightarrow{\mathbf{P}\mathbf{P}_{j_{1}}}}{\|\overrightarrow{\mathbf{P}\mathbf{P}_{j_{0}}}\times\overrightarrow{\mathbf{P}\mathbf{P}_{j_{1}}}\|},$ (15)
$\mathbf{N}(\mathbf{p})=\frac{1}{8}\sum_{j}\texttt{sign}(\mathbf{n}_{0}^{\top}\mathbf{n}_{j})\cdot\mathbf{n}_{j},$ (16)
where $\mathbf{N}\in\mathbb{R}^{3\times 1\times H\times W}$. $\mathbf{n}_{0}$ indicates the surface vector calculated with the first pair in $\mathcal{P}(\mathbf{P})$. To avoid cancellation during averaging, we align the directions of all estimated vectors with $\mathbf{n}_{0}$ by applying a sign operation on their inner products. The proposed geometric prior-guided normal consistency loss can thus be formulated as,
$\mathcal{L}_{\text{SNC}}=1-\hat{\mathbf{N}}^{\top}\mathbf{N}^{\text{DA}},$ (17)
where $\hat{\mathbf{N}}$ and $\mathbf{N}^{\text{DA}}$ indicate surface normal vector map obtained with estimated depth and scale-ambiguous depth map generated from DA, respectively. Notably, as both $\hat{\mathbf{N}}$ and $\mathbf{N}^{\text{DA}}$ are normalized to unit length, $\mathcal{L}_{\text{SNC}}$ could also be regarded as a cosine loss between these vectors. See Fig.3 for examples of surface normal vector map visualization.
Moreover, we augment this loss to $\tilde{\mathcal{L}}_{\text{SNC}}$ by employing the reconstructed spatial dense depth in previous subsection and form spatial surface normal vector map. We prove that the scale- and shift-invariance still hold for surface normal vector map generated with reconstructed spatial dense depth $\tilde{\mathbf{D}}$ and $\tilde{\mathbf{D}}^{\text{DA}}$. Detailed deduction of this can be found in Supplementary material provided along.
@@
To facilitate edge-aware estimation, in prior work of Moon et al. [15], a ground-contacting prior was introduced to mitigate erroneous depth estimation for dynamic objects by penalizing smoothness transitions between dynamic objects and ground plane, thereby encouraging alignment of the estimated depth of dynamic objects with their contacting ground points. In this work, instead of relying on segmentation cues, we leverage the output of DA not only as a geometric smoothness prior but also as an edge regularizer. This design enhances edge-aware depth estimation, promotes coherent depth transitions, and stabilizes learning in regions where photometric supervision is unreliable. Specifically, we enforce global consistency between the mean-normalized inverse depth gradients of our estimated depth and those derived from DA, which produces depth (or disparity) estimates with clear object-level silhouettes. The resulting loss function is formulated as,
@@
III-C Adaptive Joint Motion Learning
Pose estimation is a critical component for enabling pixel warping and subsequent view synthesis. Unlike prior approaches that estimate a joint motion in a fixed coordinate frame using features aggregated from all views, such as SurroundDepth [5] and VFDepth [6], or that assume a fixed camera motion with view-specific features like CVCDepth [13], we propose an adaptive joint motion learning strategy. Our approach encourages the network to learn and emphasize informative cues for structure-from-motion (SfM) learning, allowing pose estimation to adapt to varying view contributions. Technical details of the motion learning methods from previous studies are presented in Supplementary material.
Specifically, we leverage feature maps extracted from all cameras, denoted as $\boldsymbol{f}=\{\boldsymbol{f}_{i}\}_{i=1}^{N}$. We first apply spatial average pooling, followed by averaging across the camera dimension to obtain a global feature representation $\bar{\boldsymbol{f}}_{N}\in\mathbb{R}^{C}$ that aggregates holistic multi-view information. Instead of uniformly averaging camera features, we introduce a learnable fully connected network $\xi$ to predict a weight vector $\boldsymbol{\omega}$, enabling adaptive emphasis on informative views for pose estimation. The architecture of the proposed motion learning module is illustrated in Fig.2(a) and is formulated as follows.
where $\mathcal{P}_{\text{de}}$ denotes pose decoder.
III-D Geometric Feature Representation Enhancement
CLIP [10] demonstrates strong capability in capturing high-level semantic representations encoding object-level and scene-level priors, which can be complementary to low-level visual features. Motivated by this, we introduce CLIP as an auxiliary semantic encoder to enhance geometry-aware feature representations in the depth estimation pipeline. As illustrated in Fig. 2(b), CLIP model is employed to provide high-level semantic cues, which are adaptively fused with image features through a cross-modal attention mechanism, enabling the depth network to leverage semantic consistency for more robust geometric reasoning.
Specifically, we extract semantic representations using a frozen CLIP model, yielding token $\boldsymbol{c}=\texttt{CLIP}(\mathbf{I})\in\mathbb{R}^{N\times T}$, where $T$ denotes the number of semantic tokens. In parallel, the depth encoder produces multi-scale image features. To facilitate cross-modal attention, we apply a convolutional projection $g_{\phi}(\cdot)$ and a spatial resizing operation on the output of depth encoder $\boldsymbol{x}\in\mathbb{R}^{N\times C\times h\times w}$,

### v2

# GeoSurDepth: Harnessing Foundation Model for Spatial Geometry Consistency-Oriented Self-Supervised Surround-View Depth Estimation

@@ Abstract
Accurate surround-view depth estimation provides a competitive alternative to laser-based sensors and is essential for 3D scene understanding in autonomous driving. While empirical studies have proposed various approaches that primarily focus on enforcing cross-view constraints at photometric level, few explicitly exploit the rich geometric structure inherent in both monocular and surround-view setting. In this work, we propose GeoSurDepth, a framework that leverages geometry consistency as the primary cue for surround-view depth estimation. Concretely, we utilize vision foundation models as pseudo geometry priors and feature representation enhancement tool to guide the network to maintain surface normal consistency in spatial 3D space and regularize object- and texture-consistent depth estimation in 2D. In addition, we introduce a novel view synthesis pipeline where 2D-3D lifting is achieved with dense depth reconstructed via spatial warping, encouraging additional photometric supervision across temporal and spatial contexts, and compensating for the limitations of target-view image reconstruction. Finally, a newly-proposed adaptive joint motion learning strategy enables the network to adaptively emphasize informative spatial geometry cues for improved motion reasoning. Extensive experiments on KITTI, DDAD and nuScenes demonstrate that GeoSurDepth achieves SoTA performance, validating the effectiveness of our approach. Our framework highlights the importance of exploiting geometry coherence and consistency for robust self-supervised depth estimation.
@@ I Introduction
Depth estimation is a fundamental task for 3D scene understanding in autonomous driving. In recent years, self-supervised monocular depth estimation has emerged as a promising approach for 3D perception, eliminating the need for dense groundtruth annotations and making vision-based solutions attractive for large-scale, low-cost deployment [25]. By leveraging photometric reconstruction between consecutive frames or stereo pairs [18][13], these methods can learn depth directly from raw image sequences. Classical self-supervised approaches, particularly those based on monocular video, typically enforce photometric and smoothness constraints to regularize depth estimations through structure-from-motion (SfM). While effective in single-view scenarios, these methods often suffer from scale ambiguity, temporal inconsistency, and limited geometric reasoning, especially in complex scenes with dynamic objects or low texture regions.
@@
Recently, surround-view depth estimation has received growing attention in autonomous driving and robotics, where multiple cameras collectively capture a $360^{\circ}$ field of view (FoV) [8]. Accurate depth estimation in this setting is crucial for robust scene understanding, obstacle avoidance, and multi-camera fusion. However, extending self-supervised depth estimation to surround-view setups introduces additional challenges, such as cross-view consistency and spatial alignment across cameras. Although several empirical studies have begun addressing these challenges, existing methods still fall short of fully exploiting the rich geometric relationships present in overlapping views, and often fail to properly disentangle photometry, motion, and spatial geometry for consistent multi-camera depth estimation.
In this work, we propose GeoSurDepth, a framework designed to address the challenges of geometry consistency in surround-view depth estimation through simple yet effective and intuitive strategies. Our key contributions are summarized as follows: (1) We leverage the powerful vision foundation model DepthAnything (DA) V2 as an indirect pseudo-prior for spatial geometry guidance in surround-view settings. This facilitates accurate and edge-aware self-supervised depth estimation by enforcing 3D surface normal consistency and regularizing object- and texture-level consistency in 2D. Furthermore, we introduce a cross-modal attention module based on CLIP within the depth network to enhance geometric and semantic feature representation. (2) We propose a novel image synthesis approach, where dense depth reconstructed via spatial warping is utilized to achieve 2D-3D lifting, enabling photometric supervision across temporal, spatial, and spatial-temporal domains. This provides a complementary supervision signal, compensating for the limitations of image reconstruction using depth estimated in the target view only. (3) An adaptive joint motion learning strategy is introduced to enhance the network’s interpretability in emphasizing informative camera views for motion cues and learning.
In general, in this work, we aim to fully exploit geometry consistency as priors or cues to facilitate depth estimation by tailoring loss function with geometric priors and features, adapting geometry-driven motion learning and enhancing feature representation from geometric perspective.
@@ II Related Works
FSM [8] is the first work to introduce self-supervised depth estimation to the surround-view setting, aiming to achieve $360^{\circ}$ dense depth perception. By additionally incorporating photometric reconstruction losses in spatial and spatial-temporal contexts, together with a multi-camera pose consistency constraint, FSM enables metric depth estimation by explicitly exploiting spatial relationship across views. VFDepth [10] adopts a unified volumetric feature fusion strategy, enabling depth estimation from arbitrary viewpoints. In addition, it proposes a canonical motion estimation strategy that provides a global constraint for the surround-view system and derives per-camera motion via extrinsics-based motion distribution. To further constrain motion estimation and enhance cross-view interaction, SurroundDepth [15] estimates joint motion employs a Cross-View Transformer to enrich multi-view feature representations. Subsequently, MCDP [16] leverages the output of a pre-trained DepthAnything V1 model [19] as pseudo-depth for conditional denoising learning. By iteratively refining cross-view features as conditional inputs, MCDP further improves depth estimation performance. Despite these advances, existing methods primarily focus on pose consistency, feature fusion, or depth refinement, while the explicit geometric consistency and constraints of depth or motion estimation across views remains underexplored, limiting their ability to fully exploit the structural relationships inherent in surround-view camera arrays.
@@ III-A Problem Formulation
Figure 2: Network architecture of GeoSurDepth. Outputs of DA serve as surround-view geometry priors. Surround-view images at the target frame are first resized to $(518,518)$ before being fed into DA, and the output are interpolated back to original resolution. (a) Cross-modal attention mechanism: For CLIP model, input images are resized to $(214,214)$ for token extraction; (b) Adaptive joint motion learning.
We formulate surround-view depth estimation in a self-supervised manner under the conventional SfM paradigm, where dense depth and ego-motion are jointly learned from multi-camera image sequences. An overview of the proposed architecture is shown in Fig.2. The framework consists of trainable depth and pose network, together with frozen vision foundation models including DepthAnything V2 [20] and CLIP [14] model.
For network training, given surround-view images $\{\mathbf{I}_{i}^{t}\}_{i=1}^{N}$ captured by $N$ cameras at time $t$, encoder of depth network extracts multi-view features, which are jointly enhanced by fusing CLIP outputs via a cross-modal attention mechanism to improve geometric-semantic coherence. The enhanced features are then passed to depth decoder to produce surround-view depth estimates $\{\hat{\mathbf{D}}_{i}^{t}\}_{i=1}^{N}$, which would be subsequently used for 3D reconstruction and view synthesis within the SfM-based framework. To provide explicit geometric guidance, images at target time are also fed into DA, whose outputs serve as pseudo geometry priors that guide depth network toward geometry-consistent estimations.
@@ III-B Spatial Geometry Priors-Guided Self-Supervised Training
$\mathbf{p}_{ij}^{t\rightarrow t^{\prime}}=\mathbf{\Pi}_{ij}^{t\rightarrow t^{\prime}}\mathbf{p}_{i}^{t},~\tilde{\mathbf{I}}_{ij}^{t\rightarrow t^{\prime}}(\mathbf{p})=\left<\mathbf{I}_{j}^{t^{\prime}}\right>_{\mathbf{p}_{ij}^{t\rightarrow t^{\prime}}},$ (1)
@@
$\tilde{\mathbf{D}}_{j}(\mathbf{p})=\left<(\tilde{\mathbf{P}}_{j})_{z}\right>_{\mathbf{p}_{ij}},$ (8)
where $\tilde{\mathbf{P}}_{j}$ denotes a lifted 3D point from camera $j$ in coordinate frame of camera $i$. $(\cdot)_{z}$ implies $z$ value of a point cloud.
@@
$\mathcal{L}_{\text{SDC}}=\|\mathbf{D}_{i}-\tilde{\mathbf{D}}_{j}\|_{1},$ (9)
@@
Geometric prior-guided surface normal consistency (SNC) loss. Vision foundation models, trained on large-scale datasets with strong generalization capability, can provide valuable prior information for network training. MCDP [16] leverages depth estimations from a pre-trained DepthAnything V1 model as pseudo-depth for conditional denoising learning. VFM-Depth [23] incorporates DINOv2 into the depth encoder to provide universal and stable semantic features. Ninan et al. [12] used DA V2 for pseudo-supervised training. Instead of directly using depth outputs from foundation models as priors, we compute the corresponding surface normal vector maps and enforce 3D consistency in a scale- and shift-invariant manner. We present the theoretical derivation as follows.
Concretely, the output of DA is first normalized to the range $[0,1]$ to mitigate shift effects, yielding a disparity-like representation (see Fig.3). This representation is then converted into a pseudo scale-ambiguous depth map by applying a clipping range based on the estimated depth (refer to Section 1.1 of Supplementary material for more details). We exploit this to compute surface normal vector map, which is inherently invariant to scale ambiguity. Specifically, we assume $D^{\text{true}}\approx\gamma D^{\text{DA}}$, where $\gamma$ denotes an unknown scale factor. To obtain surface normal vector, we lift each pixel $\mathbf{p}$ to 3D space as $\mathbf{P}=D^{\text{true}}\mathbf{K}^{-1}\mathbf{p}$. Likewise, we lift two neighboring pixels $\mathbf{p}_{1}$ and $\mathbf{p}_{2}$, chosen such that $\cos(\overrightarrow{\mathbf{p}\mathbf{p}_{1}},\overrightarrow{\mathbf{p}\mathbf{p}_{2}})=1$. The surface normal vector at $\mathbf{P}$ can be computed via cross-product as,
@@
Specifically, we consider eight neighbors of a pixel $\mathbf{p}$ and construct four ordered pixel pairs $\mathcal{P}(\mathbf{p})=\{(\mathbf{p}_{j_{0}},\mathbf{p}_{j_{1}})\}_{j=1}^{4}$, following the strategy proposed in [17]. For each pair, the offset vectors relative to $\mathbf{p}$ are mutually perpendicular and arranged in a counterclockwise order. We lift these pixels using depth estimation by the depth network and the pseudo depth of DA, yielding the corresponding 3D point pairs $\mathcal{P}(\mathbf{P})=\{\mathbf{P}_{j_{0}},\mathbf{P}_{j_{1}}\}_{j=1}^{4}$. For each pair, we compute their cross product as described above, from which the surface normal vector map $\mathbf{N}$ is obtained as,
$\mathbf{n}_{j}(\mathbf{P})=\overrightarrow{\mathbf{P}\mathbf{P}_{j_{0}}}\times\overrightarrow{\mathbf{P}\mathbf{P}_{j_{1}}}/\|\overrightarrow{\mathbf{P}\mathbf{P}_{j_{0}}}\times\overrightarrow{\mathbf{P}\mathbf{P}_{j_{1}}}\|,$ (15)
$\mathbf{N}(\mathbf{p})=\frac{1}{4}\sum_{j}\texttt{sgn}(\mathbf{n}_{0}^{\top}\mathbf{n}_{j})\cdot\mathbf{n}_{j},$ (16)
where $\mathbf{N}\in\mathbb{R}^{3\times 1\times H\times W}$. $\mathbf{n}_{0}$ indicates the surface vector calculated with the first pair in $\mathcal{P}(\mathbf{P})$. To avoid cancellation during averaging, we align the directions of all estimated vectors with $\mathbf{n}_{0}$ by applying a sgn operation on their inner products. The proposed geometric prior-guided normal consistency loss can thus be formulated as,
$\mathcal{L}_{\text{SNC}}=1-\|\hat{\mathbf{N}}^{\top}\mathbf{N}^{\text{DA}}\|_{1},$ (17)
where $\hat{\mathbf{N}}$ and $\mathbf{N}^{\text{DA}}$ indicate surface normal vector map obtained with estimated depth and scale-ambiguous depth map of DA, respectively. Notably, as both $\hat{\mathbf{N}}$ and $\mathbf{N}^{\text{DA}}$ are normalized to unit length, $\mathcal{L}_{\text{SNC}}$ could also be regarded as a cosine loss between these vectors. Yet, we use L1 norm to remove the directional ambiguity of surface normals, making the loss invariant to sign flips, which would otherwise penalize orientation discrepancies that have no physical significance when normals are derived from pseudo depth.
Moreover, we augment this loss to $\tilde{\mathcal{L}}_{\text{SNC}}$ by employing the reconstructed spatial dense depth in previous subsection and form spatial surface normal vector map. We prove that the scale- and shift-invariance still hold for surface normal vector map generated with reconstructed spatial dense depth $\tilde{\mathbf{D}}$ and $\tilde{\mathbf{D}}^{\text{DA}}$. Detailed deduction of this can be found in Section 1.3 of Supplementary material provided along.
@@
To facilitate edge-aware estimation, in prior work of Moon et al. [11], object masks introduced to mitigate erroneous depth estimation for dynamic objects by penalizing smoothness transitions between dynamic objects and ground plane, thereby encouraging alignment of the estimated depth of dynamic objects with their contacting ground points. In this work, instead of relying on segmentation cues, we leverage the output of DA not only as a geometric smoothness prior but also as an edge regularizer. This design enhances edge-aware depth estimation, promotes coherent depth transitions, and stabilizes learning in regions where photometric supervision is unreliable. Specifically, we enforce global consistency between the mean-normalized inverse depth gradients of our estimated depth and those derived from DA, which produces depth (or disparity) estimates with clear object-level silhouettes. The resulting loss function is formulated as,
@@
In summary, we use surface normal vector map and disparity representation of DA as pseudo geometry priors only. In Fig.4 we present an imperfect example of DA output.
Figure 4: Indication of pseudo geometric priors by DepthAnything V2: Depth of a truck in the front is wrongly estimated.
@@
III-C Geometric Feature Representation Enhancement
CLIP demonstrates strong capability in capturing high-level semantic representations encoding object- and scene-level priors, which can be complementary to low-level visual features. Motivated by this, we introduce CLIP as an auxiliary semantic encoder to enhance geometry-aware feature representations in the depth estimation pipeline. As illustrated in Fig.2(b), CLIP model is employed to provide high-level semantic cues, which are adaptively fused with image features through a cross-modal attention mechanism, enabling the depth network to leverage semantic consistency for more robust geometric reasoning.
Specifically, we extract semantic representations using a frozen CLIP model, yielding token $\boldsymbol{c}=\texttt{CLIP}(\mathbf{I})\in\mathbb{R}^{N\times T}$, where $T$ denotes the number of semantic tokens. In parallel, the depth encoder produces multi-scale image features. To facilitate cross-modal attention, we apply a convolutional projection $g_{\phi}(\cdot)$ and a spatial resizing operation on the last layer output of depth encoder $\boldsymbol{x}\in\mathbb{R}^{N\times C\times h\times w}$,